# RemoteReasoner: Towards Unifying Geospatial Reasoning Workflow

### v1

@@ Abstract
Remote sensing imagery presents vast, inherently unstructured spatial data, demanding sophisticated reasoning to interpret complex user intents and contextual relationships beyond simple recognition tasks. In this paper, we aim to construct an Earth observation workflow to handle complex queries by reasoning about spatial context and user intent. As a reasoning workflow, it should be somewhat autonomous, where predefined ground-truth reasoning paths do not constrain the learning process. Furthermore, its architecture ought to be unified yet flexible, enabling the model to perform diverse reasoning tasks with distinct output formats through a single forward pass. Existing remote sensing approaches fail to address these requirements, as they rely on supervised fine-tuning paradigms that constrain the autonomy of reasoning. To this end, we propose RemoteReasoner, a flexible and robust workflow for remote sensing reasoning tasks. The design of RemoteReasoner integrates a multi-modal large language model (MLLM) for interpreting user instructions and localizing targets, together with task adaptation strategies that enable multi-granularity output generation. In contrast to existing methods, our framework is trained with reinforcement learning (RL) to endow the MLLM sufficient autonomy for precise reasoning. At the inference stage, our adaptation strategies enable diverse output formats at inference time without requiring task-specific decoders or further fine-tuning. Preliminary experiments demonstrated that RemoteReasoner achieves remarkable performance across multi-granularity reasoning tasks, including region-level and pixel-level. Additionally, our framework enables novel capabilities such as the contour extraction task beyond the reach of existing reasoning pipelines.
@@ 1 Introduction
Remote sensing analysis has evolved beyond simple object detection Zou et al. (2023) or classification Adegun et al. (2023), demanding sophisticated reasoning capabilities to unlock its full potential for complex, real-world decision-making. Although traditional pixel- Yan et al. (2023) or object-centric Liu et al. (2025) approaches are effective for narrow tasks, they fundamentally lack the ability to interpret nuanced user intents, infer implicit spatial relationships, or dynamically chain logical steps to answer intricate queries. Reasoning Jaech et al. (2024), Xu et al. (2025) bridges this gap by enabling systems to dynamically decipher the underlying goals of ambiguous or high-level instructions (e.g., "Find areas at risk of flooding near critical infrastructure after heavy rain").
@@
Numerous efforts Lai et al. (2024), Ren et al. (2024), Chen et al. (2024b) have explored reasoning tasks in the natural image domain. For instance, LISA Lai et al. (2024) utilizes multi-modal large language models (MLLMs) Carolan et al. (2024) for reasoning-driven segmentation tasks. However, these approaches rely on supervised fine-tuning (SFT) paradigms to train the models. Such frameworks constrain model autonomy, ultimately compromising generalization capabilities. This situation can be manifested in substantial performance degradation on out-of-distribution samples and catastrophic forgetting of general capabilities. Similarly, the remote sensing community has seen pioneering works Li et al. (2025), Yao et al. (2025b). SegEarth-R1 Li et al. (2025) introduces a geospatial pixel reasoning task that integrates a hierarchical MLLM and a mask decoder to generate both reasoning processes and segmentation outputs. However, it remains constrained by SFT paradigms and exclusively supports segmentation tasks.
@@
These limitations motivate us to design a reasoning workflow that autonomously interprets user intent through self-directed thinking and efficiently generates geospatial interpretations in diverse output formats tailored for downstream intelligence analysis. Firstly, the requirement for self-thinking should be fulfilled through an open-ended learning paradigm. This aligns naturally with the emerging reinforcement learning (RL) framework Guo et al. (2025), Shen et al. (2025). Rather than SFT approaches that demand strictly annotated reasoning processes, RL explores the self-evolution potential of models to cultivate reasoning capabilities, thereby naturally enhancing model generalization. Secondly, to support flexible multi-task adaptation, we should avoid redundant training through task transformation at the decision-making level.
@@
Followed by these insights, we propose RemoteReasoner, a reasoning workflow that integrates a robust MLLM and a flexible inference strategy. Building upon the pixel-level reasoning task proposed in SegEarth-R1 Li et al. (2025), RemoteReasoner further implements two novel tasks through autonomous reasoning: 1) region-level: geospatial region reasoning, and 2) object-level: geospatial contour reasoning. These three distinct tasks can be achieved through a single MLLM inference followed by a unified decision-level task transformation pipeline, significantly improving output efficiency for multi-granularity reasoning tasks.
@@
Extensive experimental results demonstrate that RemoteReasoner achieves state-of-the-art performance on pixel-level and region-level reasoning tasks, with enhanced autonomous reasoning capabilities and superior generalization. For instance, RemoteReasoner achieves over 40% accuracy improvement in region-level geospatial reasoning tasks. Moreover, compared to existing methods, RemoteReasoner supports a broader range of tasks without introducing additional decoders.
@@
We propose RemoteResoner, a unified yet flexible reasoning workflow for Earth observation, which reveals that integrating pure RL and task transformation strategies could handle multi-granularity reasoning tasks.
@@ 3 RemoteReasoner
In this section, we introduce our method of building a workflow for multi-level geospatial reasoning tasks. The overall framework is represented in Fig. 1.
@@ 3.2 Reasoning Model
We utilize Qwen2.5-VL-7B Bai et al. (2025) as our reasoning model . Despite Qwen2.5-VL’s superior performance in general-domain applications, direct adaptation to remote sensing imagery introduces domain knowledge discrepancy. Moreover, its reasoning capabilities require significant enhancement to address complex geospatial tasks. Therefore, we require fine-tuning it through reinforcement learning on remote sensing data.
@@ 3.4 Inference Workflow
Geospatial Pixel Reasoning: The bounding box serves as a prompt for the Segment Anything Model (SAM) . SAM generates a binary mask isolating the target:
@@
Geospatial Contour Reasoning: The target contour is derived from via morphological boundary detection:
@@ 4.4.1 Can RemoteReasoner Do Other Tasks?
We evaluate RemoteReasoner on Visual Grounding and Referring Expression Segmentation tasks, with results summarized in Table X. Comparative experiments are conducted against state-of-the-art LLM-based approaches, including general-domain methods (MiniGPT-V2 Zhu et al. (2023), LISA Lai et al. (2024), PixelLM Ren et al. (2024)) and remote sensing-specific frameworks (Falcon Yao et al. (2025a), LHRS-Bot Muhtar et al. (2024), GeoGround Zhou et al. (2024)). Notably, all remote sensing methods were trained on task-specific data, while RemoteReasoner’s zero-shot performance achieves 80% of their supervised results. It demonstrates our framework’s strong generalization capabilities with reinforcement learning.
4.4.2 Can RemoteReasoner Recognize Unseen Category?
@@
5 Conclusion
@@
We present RemoteReasoner, a novel workflow advancing autonomous geospatial reasoning. Overcoming limitations of SFT-based approaches, it leverages RL strategy to cultivate robust reasoning capabilities while preserving MLLM generalization. Its core innovation is a unified inference pipeline enabling efficient multi-granularity task transformation (pixel-, region-, and object-level) from a single MLLM output, eliminating redundant per-task decoders. RemoteReasoner achieves SOTA results, including 40% accuracy gains in region-level reasoning, and demonstrates superior generalization across diverse, complex queries. It provides a flexible foundation for downstream geospatial intelligence requiring nuanced intent interpretation and multi-format outputs. This work is still in progress, and the model and code will be open-sourced in the next version.

### v2

@@ Abstract
Remote sensing imagery presents vast, inherently unstructured spatial data, necessitating sophisticated reasoning to interpret complex user intents and contextual relationships beyond simple recognition tasks. In this paper, we aim to construct an Earth observation workflow to handle complex queries by reasoning about spatial context and user intent. As a reasoning workflow, it should autonomously explore and construct its own inference paths, rather than being confined to predefined ground-truth sequences. Ideally, its architecture ought to be unified yet generalized, possessing capabilities to perform diverse reasoning tasks through one model without requiring additional fine-tuning. Existing remote sensing approaches rely on supervised fine-tuning paradigms and task-specific heads, limiting both autonomous reasoning and unified generalization. To this end, we propose RemoteReasoner, a unified workflow for geospatial reasoning. The design of RemoteReasoner integrates a multi-modal large language model (MLLM) for interpreting user instructions and localizing targets, together with task transformation strategies that enable multi-granularity tasks, including object-, region-, and pixel-level. In contrast to existing methods, our framework is trained with reinforcement learning (RL) to endow the MLLM sufficient reasoning autonomy. At the inference stage, our transformation strategies enable diverse task output formats without requiring task-specific decoders or further fine-tuning. Experiments demonstrated that RemoteReasoner achieves state-of-the-art (SOTA) performance across multi-granularity reasoning tasks. Furthermore, it retains the MLLM’s inherent generalization capability, demonstrating robust performance on unseen tasks and out-of-distribution categories.
@@ 1 Introduction
Remote sensing analysis has evolved beyond simple object detection Zou et al. (2023), Liu et al. (2025a) or classification Adegun et al. (2023), increasingly requiring stronger capabilities to support complex real-world decision-making. Although existing approaches Yao et al. (2025b), Liu et al. (2024a) are effective for multiple tasks, they struggle to interpret nuanced user intent and infer implicit spatial relationships for intricate queries. Reasoning Jaech et al. (2024), Xu et al. (2025) bridges this gap by enabling systems to dynamically decipher the underlying goals of ambiguous or high-level instructions (e.g., “Find areas at risk of flooding near critical infrastructure after heavy rain"). Such a geospatial workflow would conveniently accommodate varied geoscience applications Zhao et al. (2024).
@@
Numerous efforts Lai et al. (2024), Ren et al. (2024), Chen et al. (2024b) have explored reasoning tasks in the natural image domain. For instance, LISA Lai et al. (2024) utilizes multi-modal large language models (MLLMs) Carolan et al. (2024) for reasoning-driven segmentation tasks. However, these approaches rely on supervised fine-tuning (SFT) paradigms to train the models. Such frameworks constrain model autonomy, ultimately compromising generalization capabilities. This situation can be manifested in substantial performance degradation on out-of-distribution samples and catastrophic forgetting of general capabilities Liu et al. (2025b). Similarly, the remote sensing community has seen pioneering works Li et al. (2025), Zhang et al. (2023), Wang et al. (2024). SegEarth-R1 Li et al. (2025) introduces a geospatial pixel reasoning task that integrates a hierarchical MLLM and a mask decoder to generate both reasoning processes and segmentation outputs. However, it remains constrained by SFT paradigms and exclusively supports segmentation tasks.
@@
These limitations motivate us to design a reasoning workflow that autonomously interprets user intent through self-directed thinking. It should also efficiently generate geospatial interpretations in diverse output formats, which are tailored for various downstream intelligence analysis. Firstly, the requirement for self-thinking should be fulfilled through an open-ended learning paradigm. This aligns naturally with the emerging reinforcement learning (RL) frameworks Guo et al. (2025), Shen et al. (2025). Rather than SFT approaches that demand strictly annotated reasoning processes, RL explores the model’s self-evolution potential to cultivate reasoning capabilities. Secondly, to flexibly support multi-granularity tasks, we should avoid redundant task decoders and training through decision-level task transformation.
@@
Followed by these insights, we propose RemoteReasoner, a reasoning workflow that integrates a robust MLLM and a flexible inference strategy. As shown in Fig. 1, building upon the pixel-level reasoning task proposed in SegEarth-R1 Li et al. (2025), RemoteReasoner further implements two novel tasks through autonomous reasoning: 1) region-level: geospatial region reasoning, and 2) object-level: geospatial contour reasoning. These three distinct tasks can be achieved through a single MLLM inference followed by a unified decision-level task transformation pipeline, significantly improving output efficiency for multi-granularity reasoning tasks.
@@
Extensive experimental results demonstrate that RemoteReasoner achieves state-of-the-art performance on multi-granularity reasoning tasks, evidencing its enhanced autonomous reasoning capabilities and superior generalization. For instance, RemoteReasoner achieves 3.67% accuracy improvement over SegEarth-R1 in pixel-level geospatial reasoning tasks. On unseen tasks (e.g., , Referring Segmentation), our framework significantly outperforms competing models (e.g., , PixelLM, LISA) in performance.
@@
We propose RemoteResoner, a unified yet generalized reasoning workflow for Earth observation, which reveals that integrating pure RL and task transformation strategies could handle multi-granularity reasoning tasks.
@@ 3 RemoteReasoner
In this section, we introduce our method of building a workflow for multi-granularity geospatial reasoning tasks. The overall framework is represented in Fig. 2.
@@ 3.2 Reasoning Model
We utilize Qwen2.5-VL-7B-instruct Bai et al. (2025) as our reasoning model . Despite Qwen2.5-VL’s superior performance in general-domain applications, direct adaptation to remote sensing imagery introduces domain knowledge discrepancy. Moreover, its reasoning capabilities require significant enhancement to address complex geospatial tasks. Therefore, we require fine-tuning it through reinforcement learning on remote sensing data.
@@ 3.4 Inference Workflow
Geospatial Pixel Reasoning: The bounding box and its center points serve as prompts for the Segment Anything Model (SAM) . SAM generates a binary mask isolating the target:
@@
Geospatial Contour Reasoning: The target contour is derived via morphological boundary detection:
@@ 4.3.3 Geospatial Contour Reasoning
| Output Format | Training Method | ||
| Text2Mask | SFT | 22.01 | 22.97 |
| RL | 18.74 | 19.21 | |
| Bbox2Mask | SFT | 60.25 | 62.19 |
| RL | 69.13 | 70.96 |
@@ 4.4.1 Can RemoteReasoner Do Other Tasks?
We evaluate RemoteReasoner on Visual Grounding and Referring Expression Segmentation tasks, with results summarized in Tab. 4. Comparative experiments are conducted against state-of-the-art LLM-based approaches, including general-domain methods (MiniGPT-V2 Zhu et al. (2023), LISA Lai et al. (2024), PixelLM Ren et al. (2024)) and remote sensing-specific frameworks (Falcon Yao et al. (2025a), LHRS-Bot Muhtar et al. (2024), GeoGround Zhou et al. (2024)). Notably, all remote sensing methods were trained on task-specific data, while RemoteReasoner’s zero-shot performance achieves 80% of their supervised results. Under identical zero-shot settings, our framework achieves gains of 4.12% on Visual Grounding and 19.32% on Referring Expression Segmentation tasks.
Furthermore, as the workflow’s core component is a multi-modal large language model, we verify preservation of its inherent capabilities through multi-turn dialogue evaluation (Fig. 4). The model first performs image captioning, then accurately answers content-based questions about the image. Additional results are provided in the supplementary material. These results demonstrate our framework’s strong generalization capabilities with RL.
4.4.2 Can RemoteReasoner Recognize Other Category?
@@
4.4.3 Qualitative Analysis
The qualitative results of the three supported tasks in our workflow are presented in Fig. 5. In the first sample row, the user seeks a venue for group sports, while a baseball field nearly blends into the background in the image. RemoteReasoner correctly reasons and locates the target, simultaneously yielding its segmentation mask and contour. The second row shows a relatively blurred image of a medical helipad. Based on its function of providing emergency rescue services, RemoteReasoner accurately locates and segments the target area along with its contour. These visualizations indicate that our method can adapt to challenging reasoning tasks.
4.5 Ablation Studies
4.5.1 Different Pixel-level Output Formats
Recently, RemoteSAM Yao et al. (2025b) validates that segmentation, as a fundamental visual output unit, enables bottom-up transformation to other vision tasks. Therefore, if LLMs could directly generate masks in textual representations, the architectural design holds promise for greater unification. To explore this hypothesis, we conducted experiments using such formulations, as summarized in Tab. 6. However, textual outputs prove ill-suited for this predictive task, with both SFT and RL strategies exhibiting significant performance degradation compared to visual-based approaches. In contrast, directly outputting bounding boxes and then generating masks through SAM significantly outperforms textual mask representations (22.01 vs. 60.25). When incorporating RL strategies, the accuracy further improves from 60.25 to 69.13. Based on these findings, we adopt the Mask2Contour approach in our workflow design.
4.5.2 Data Scaling Experiment
To validate data impact on model performance, we evaluate it by training with 20%, 50%, and 100% subsets of the full training data. As illustrated in Fig. 6 (a), the results demonstrate that model performance improves with increasing data scale. The full dataset (100%) achieves the best performance, indicating the current data scale contains no redundancy. These findings suggest that geospatial reasoning tasks retain untapped potential that can be unlocked through larger data scales.
4.5.3 Inference Cost Ablation
We further evaluate inference speed across different contour extraction strategies, as shown in Fig. 6 (b). The Mask2Contour approach directly applies morphological operations without introducing additional parameters, achieving significantly faster processing speed (0.09s per image). In contrast, EPOC, as a deep learning model with 84.6MB parameters, demonstrates inferior inference speed (0.43s per image). This comparison highlights the efficiency advantages of our proposed workflow.
5 Conclusion & Limitations
@@
Conclusion. We present RemoteReasoner, a novel workflow advancing autonomous geospatial reasoning. Overcoming limitations of SFT-based approaches, it leverages RL strategy to cultivate robust reasoning capabilities while preserving MLLM generalization. Its core innovation is a unified inference pipeline enabling efficient multi-granularity task transformation (pixel-, region-, and object-level) from a single MLLM output, eliminating redundant per-task decoders. RemoteReasoner achieves SOTA results, including 15% and 30% accuracy gains in region- and object-level reasoning, and demonstrates superior generalization across diverse, complex queries. It provides a flexible foundation for downstream geospatial intelligence requiring nuanced intent interpretation and multi-format outputs.
Limitations. Despite the advantages, several limitations exist for future improvement. First, model performance remains limited by current reasoning dataset scale. Progress will require more comprehensive datasets. Second, the performance on small objects needs enhancement via specialized data and architectural refinements. Additionally, exploring more tasks could advance unified geospatial reasoning.
Appendix of RemoteReasoner
A. More Implementation Details
The training pipeline leverages MS-SWIFT framework with Group Relative Preference Optimization (GRPO). Experiments use 8 NVIDIA L20 GPUs with DeepSpeed Zero-3 for memory optimization. We initialize weights from Qwen2.5-VL-7B-Instruct and apply LoRA adaptation (rank 8, ) to all linear layers. The training configuration employs a global batch size of 512 (8 per-device batches 8 gradient accumulation steps 8 GPUs), bfloat16 precision, and flash attention for computational efficiency. We train for 24 epochs with a learning rate of , linear warmup (5% of steps), and gradient checkpointing. Input sequences are capped at 4096 tokens with image resolution limited to 802816 pixels. The optimizer uses AdamW with weight decay disabled. For RLHF alignment, we utilize a temperature of 0.9 during generation.
B. More Experiments
B.1 Balance Factors Sensitive Analysis
| 0.3 | 0.5 | 0.7 | 1 | |
| 0.3 | 47.24 | 46.73 | 46.22 | 48.01 |
| 0.5 | 45.21 | 46.85 | 48.03 | 49.26 |
| 0.7 | 47.07 | 46.90 | 49.88 | 49.05 |
| 1 | 47.78 | 49.67 | 48.95 | 50.11 |
Before training, to obtain an optimal set of balance factors for the reward function, we sampled a small portion of the dataset and trained the model using different weighting schemes for the accuracy rewards. Tab. 7 presents a sensitivity analysis of the reward function balance factors, each ranging from 0.3 to 1.0. The values represent model performance under each configuration on the validation set. Among all tested combinations, the setting where both balance factors are set to 1.0 achieves the best performance, reaching 50.11%. Compared to lower-weighted settings, the (1.0, 1.0) configuration provides stronger and more consistent supervision, leading to more effective optimization. Therefore, we adopt this configuration in our final model training.
B.2 Point Prompt Effectiveness Ablations
| Bbox | Point | ||
| 43.09 | 42.41 | ||
| 69.13 | 70.96 |
To evaluate the effectiveness of different prompt types in guiding the SAM model, we conducted a controlled experiment using bounding box (bbox) and center point prompts. Specifically, we compare two settings: using only bbox prompts, and using both bbox and point prompts, where the point is simply computed as the center of the bbox—without relying on additional points generated by the LLM. As shown in the Tab. 8, using only the bbox yields moderate performance with 43.09 and 42.41 . However, when the center point is additionally provided as a prompt, performance improves significantly to 69.13 and 70.96 . This substantial gain highlights the complementary role of point prompts in guiding the segmentation process. The result suggests that even simple, derived point inputs (like bbox centers) can substantially enhance SAM’s segmentation accuracy when used in conjunction with bounding boxes.
C. More Qualitative Results
In this section, we present additional visualizations for multi-granularity reasoning tasks. Results for region-level and pixel-level reasoning are illustrated in Fig. 7, while contour reasoning outcomes appear in Fig. 8. Results incorporating the reasoning process are depicted in Fig. 9 and 10. Additionally, we provide image-level task results—including Visual Question Answering (VQA) and image captioning—as summarized in Fig. 11 and 12.
D Failure Case Analysis
Fig. 13 presents several illustrative failure cases of RemoteReasoner. First row: The query requests an entity that circumvents the urban core (expected: an overpass). However, the model incorrectly classifies the entire intersection as fulfilling this condition, consequently outputting its complete structure. Second row: The query requires locating the grandstand. While the model outputs the entire baseball stadium, the reasoning process leading to this output is logically sound. It is noteworthy that the ground truth annotation for this sample is also incorrect, as it fails to label the grandstand location. Third row: The model mistakenly incorporates the shadow region as part of the target object. These failure cases highlight significant areas for improvement in our method.
E. Prompts Setting
We present the prompt templates designed during dataset construction in Tab. 9.